A MICROWAVE STUDY OF PARTIALLY BOUND H_{2}O-BF_{3}

\underline{M.E. OTT}, D.L. FIACCO, S.R. WEERS, T. HOWE, Department of Chemistry, University of Minnesota, 207 Pleasant St., SE, Minneapolis, MN 55455; S.W. REEVE, Department of Chemistry and Biochemistry, Arkansas State University, PO Drawer 419, State University, AR 72467-0419; C.J. CRAMER, and K.R. LEOPOLD, Department of Chemistry, University of Minnesota, 207 Pleasant St., SE, Minneapolis, MN 55455.

We report the microwave spectrum and structure of the partially bonded donor - acceptor adduct H_{2}O-BF_{3}. H_{2}O and BF_{3} react to form two compounds, a monohydrate and a dihydrate, both with well defined stoichiometries and melting points. For the monohydrate, previous X-ray diffraction work^{a} indicates a B-O bond distance of 1.532 Å (equal to the sum of covalent radii) and a tetrahedral geometry at the boron. In the gas, however, we find a bond length of 1.90 Å and an out-of-plane distortion of the BF_{3} of only 8^{o}. Thus, while the B-O bond in the solid is an ”ordinary” dative linkage, the bond in the gas phase system appears to be only partially developed. The large changes in bond length and bond angle which occur upon crystallization (0.4 Å and 9^{o}, respectively) will be discussed in comparison with other partially bound systems which have been investigated in our laboratory. Ab initio calculations place the gas phase binding energy at 9.6 kcal/mol, which is a very reasonable value for a partially formed dative bond.

aVon D. Mootz and M. Steffen Z. Anorg. Allg. Chem. \underline{483},171,1981.